Calculate the concentrations of all species in a 0.100 M H3PO4 solution.(Ka1 = 7.1 x 10^-3

, Ka2 = 6.3 x 10^-8 and

Ka3 = 4.5 x 10^-13)


1
Expert's answer
2020-12-03T13:50:48-0500

First Dissociation

H3PO4(aq)= H+(aq) + H2PO-4(aq)

I          0.100        0         0

C         -x             +x          +x

E         (0.1-x)        x             x

Ka1= [H+] [H2PO-4]/[H3PO4]

x2/(0.1-x) =7.1× 10-3

x2/0.1=7.1×10-3

Assuming that x˂˂0.1

x2=0.1×7.1×10-3

x=0.027M

Second Dissociation

H2PO-4(aq)=H+(aq) + HPO2-4(aq)

I          0.027       0.027       0

C         -y              +y             +y

E        ( 0.027-y)   (0.027+y)   y

Ka2= [H+] [HPO2-4]/[H2PO-4] = y(0.19+y)/(0.19-y)

    y=6.3×10-8M

Assuming that y˂˂0.027

Third Dissociation

HPO2-4(aq)= H+(aq)  +     PO3-4(aq)

I          6.3×10-8     0.19          0

C         -z                 +z               +z

E          (6.3×10-8)-z  (0.027+z)  z

Ka3=[H+] [PO3-4]/[HPO2-4]

   (0.027+z) z/ (6.3×10-8)-z = 4.5×10-13

0.027z/ (6.3×10-8) = 4.5×10-13

Assuming that [H+] from third dissociation completely negligible

Z= 1.05×10-18 M